Solve Problem Using Linear Systems
13. One type of granola has 30% nuts, by mass. A second type of granola has 15% nuts. What mass of each type needs to be mixed to make 600 g of granola that will have 21% nuts?
Can you please help me with this question thank you very much!!!

Answer by edjones(8007)

Let x be grams of 30% nuts granola
.3x+.15(600-x)=600*.21
.3x+90-.15x=126
.15x=36
x=240g
600-x=360g of 15% nut granola
.
Ed